OpenStudy (anonymous):

A volume of gas, initially at P = 100,000 N/m^2 and V = 200 m^3 is heated at constant P from 300 to 330 K. How much work is done by the gas as it expands?

OpenStudy (anonymous):

I got each question on the sample exam correct except this one. Not sure how to attack it.

OpenStudy (anonymous):

Work is defined as: W = -P dV However, we don't have a final volume. So let's find it. Use the ideal gas law. PiVi=NKTi & PfVf=NKTf If we divide everything to one side, we get two equations equal to 1. We can set them equal. The ideal gas is still an ideal gas after the process, so the law has to hold in both cases. PiVi/NKTi = PfVf/NKTf --> NK cancel, P cancels b/c it's const. Vi/Ti = Vf/Tf --> Solve accordingly and find Vf = 220 m^3 Now work we can find: W = -P dV = -100000 (220 - 200) = -2000000 J